Charge is spread uniformly over the surface of a sphere of radius R. The potential at the sphere's center is V.

Find an expression for the net charge Q on the sphere.

Express your answer in terms of the variables R, V, and the Coulomb's constant k.
Answer #1

charge is spread unifromly over the surface of sphere.

hence it is a spherical shell of charge.

due to spherical shell,

potential at centre , V = kq/R

q = VR/k

where k = 1 / 4 pi e0


or q = 4 pi e0 V R